# Evaluating Robustness of Vision Foundation Models for Semantic Segmentation

### v1

@@ Abstract
In this report, we present our solution for Track 1 of the 2024 BRAVO Challenge, where a model is trained on Cityscapes and its robustness is evaluated on several out-of-distribution datasets. Our solution leverages the powerful representations learned by vision foundation models, by attaching a simple segmentation decoder to DINOv2 and fine-tuning the entire model. This approach outperforms more complex existing approaches, and achieves 1st place in the challenge. Our code is publicly available [^1] https://github.com/tue-mps/benchmark-vfm-ss.
@@ 1 Introduction
The 2024 BRAVO Challenge “aims to benchmark semantic segmentation models on urban scenes undergoing diverse forms of natural degradation and realistic-looking synthetic corruption" [17]. In this report, we present our solution for Track 1 of this challenge, where a model is trained on a single labeled semantic segmentation dataset, Cityscapes [4], and evaluated on a range of other datasets with out-of-distribution image conditions and semantic content, to evaluate the robustness of the model.
@@
In this report, we present our simple meta-approach, and evaluate several configurations. In our default configuration, we use a DINOv2 [12] VFM, attach a simple linear decoder for segmentation, and fine-tune the entire model. We assess the impact of using different model sizes, patch sizes, pre-training strategies and segmentation decoders. With our simple approach, we significantly outperform existing specialist models and obtain the 1st place in the challenge. Moreover, we make some new observations which could be of interest to future work.
@@ 2 Method
Our primary solution, which achieves 1st place in the challenge, uses the DINOv2 VFM [12]. We selected DINOv2 due to its demonstrated effectiveness in domain-generalized semantic segmentation for urban scenes [10, 7]. DINOv2, built upon the Vision Transformer (ViT) architecture [6], is pre-trained using self-supervised learning on a vast, curated dataset. We conduct experiments using all available sizes of DINOv2.
@@ 3.2 Evaluation metrics
The 2024 BRAVO Challenge evaluates methods based on a variety of metrics to assess their performance in semantic segmentation and out-of-distribution (OOD) detection. The metrics are grouped into three categories: semantic metrics, OOD metrics, and summary metrics. The semantic metrics evaluate the overall quality of the semantic segmentation predictions, while the OOD metrics assess the model’s ability to detect OOD objects. The summary metrics combine the semantic and OOD metrics to provide an overall ranking of the methods.
@@
The OOD metrics are computed on the SMIYC and Synobjs subsets for invalid pixels only. Invalid pixels in these datasets are those that are obscured by OOD objects. The OOD metrics include:
@@ 4 Results
DINOv2, ViT-L, 8x8 patch size, linear decoder 76.7 40.0 99.4 91.4 2.1 38.2
DINOv2, ViT-L, 16x16 patch size, linear decoder 75.9 41.2 99.5 92.3 1.8 37.2
DINOv2, ViT-g, 16x16 patch size, linear decoder 77.6 39.3 99.5 92.3 1.8 37.1
DINOv2, ViT-B, 16x16 patch size, linear decoder 71.7 43.3 99.4 92.3 2.1 39.5
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.5 45.1 99.2 91.8 2.5 43.8
PixOOD YOLO (Model Selection) 43.6 57.3 93.5 83.5 15.5 55.7
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 78.2 23.2 99.2 87.9 5.1 61.9
Model selection 73.6 44.8 99.1 91.8 10.0 38.3
PixOOD w/ ResNet-101 DeepLab 43.2 58.5 93.5 84.0 15.5 54.4
Ensemble C 73.9 47.4 99.1 92.5 52.7 33.7
Ensemble A 73.8 47.0 99.1 92.6 38.6 35.2
PixOOD w/ DeepLab Decoder 69.3 23.5 97.4 76.0 13.8 70.1
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 44.9 54.7 97.9 89.2 1.8 52.8
PixOOD 64.8 22.4 96.3 73.1 18.1 76.5
SegFormer-B5 67.4 24.1 97.2 77.2 30.3 71.5
ObsNet-R101-DLv3plus 65.3 32.1 98.5 87.8 45.9 62.8
Mask2Former-SwinB 67.2 13.2 90.4 47.0 55.2 82.3
Physically Feasible Semantic Segmentation 67.4 42.2 98.7 89.4 4.8 47.0
@@
Another noteworthy observation is that all our models have relatively low ECE values, indicating that they are well-calibrated, even though no explicit calibration techniques were applied. Even the DeiT-III-based model, which scores low on the overall BRAVO score, achieves a low ECE value of 1.8. Therefore, further investigation is needed to understand why the ECE values are so low.
@@ 5 Conclusion
This report presents the winning solution for Track 1 of the 2024 BRAVO Challenge. Our approach is based on fine-tuning Vision Foundation Models (VFMs) for semantic segmentation, leveraging the strong representations learned during pre-training. The results demonstrate that, just through better pre-training, a fine-tuned VFM is more robust under distribution shifts than complex specialist models, and can both accurately classify pixels and reliably estimate its confidence in these predictions. We evaluated our approach in several configurations, which has lead to several new observations. While we identify potential causes of these observations, future work is necessary to explore this further. Particularly, we believe it would be interesting to investigate (a) the cause of the low calibration error of our ViT-based models, and (b) the reason for the difference in semantic and OOD metrics with the Mask2Former decoder compared to the linear decoder. Additionally, future works could explore whether some of the more specialized out-of-distribution detection or calibration methods are still effective when used in combination with VFMs.

### v2

# First Place Solution to the ECCV 2024 BRAVO Challenge: Evaluating Robustness of Vision Foundation Models for Semantic Segmentation

@@ Abstract
In this report, we present the first place solution to the ECCV 2024 BRAVO Challenge, where a model is trained on Cityscapes and its robustness is evaluated on several out-of-distribution datasets. Our solution leverages the powerful representations learned by vision foundation models, by attaching a simple segmentation decoder to DINOv2 and fine-tuning the entire model. This approach outperforms more complex existing approaches, and achieves first place in the challenge. Our code is publicly available [^1] https://github.com/tue-mps/benchmark-vfm-ss.
@@ 1 Introduction
The ECCV 2024 BRAVO Challenge “aims to benchmark semantic segmentation models on urban scenes undergoing diverse forms of natural degradation and realistic-looking synthetic corruption" [17]. In this report, we present our solution for Track 1 of this challenge, where a model is trained on a single labeled semantic segmentation dataset, Cityscapes [4], and evaluated on a range of other datasets with out-of-distribution image conditions and semantic content, to evaluate the robustness of the model.
@@
In this report, we present our simple meta-approach, and evaluate several configurations. In our default configuration, we use a DINOv2 [12] VFM, attach a simple linear decoder for segmentation, and fine-tune the entire model. We assess the impact of using different model sizes, patch sizes, pre-training strategies and segmentation decoders. With our simple approach, we significantly outperform existing specialist models and obtain the first place in the challenge. Moreover, we make some new observations which could be of interest to future work.
@@ 2 Method
Our primary solution, which achieves first place in the challenge, uses the DINOv2 VFM [12]. We selected DINOv2 due to its demonstrated effectiveness in domain-generalized semantic segmentation for urban scenes [10, 7]. DINOv2, built upon the Vision Transformer (ViT) architecture [6], is pre-trained using self-supervised learning on a vast, curated dataset. We conduct experiments using all available sizes of DINOv2.
@@ 3.2 Evaluation metrics
The ECCV 2024 BRAVO Challenge evaluates methods based on a variety of metrics to assess their performance in semantic segmentation and out-of-distribution (OOD) detection. The metrics are grouped into three categories: semantic metrics, OOD metrics, and summary metrics. The semantic metrics evaluate the overall quality of the semantic segmentation predictions, while the OOD metrics assess the model’s ability to detect OOD objects. The summary metrics combine the semantic and OOD metrics to provide an overall ranking of the methods.
@@
The OOD metrics are computed on the SMIYC and Synobjs subsets only. Invalid pixels in these datasets are those that are obscured by OOD objects. The OOD metrics quantify whether the model attributes, as expected, less confidence to the invalid pixels. The OOD metrics include:
@@ 4 Results
DINOv2, ViT-L, 8x8 patch size, linear decoder 76.7 40.0 99.4 91.4 2.0 38.8
DINOv2, ViT-L, 16x16 patch size, linear decoder 75.9 41.2 99.5 92.3 1.7 37.8
DINOv2, ViT-g, 16x16 patch size, linear decoder 77.6 39.3 99.5 92.3 1.8 37.6
DINOv2, ViT-B, 16x16 patch size, linear decoder 71.7 43.3 99.4 92.3 2.1 40.3
DINOv2, ViT-S, 16x16 patch size, linear decoder 66.5 45.1 99.2 91.8 2.5 44.3
PixOOD YOLO (Model Selection) 43.6 57.3 93.5 83.5 15.1 55.8
DINOv2, ViT-g, 16x16 patch size, Mask2Former decoder 78.2 23.2 99.2 87.9 5.0 63.6
Model selection 73.6 44.8 99.1 91.8 9.2 40.2
PixOOD w/ ResNet-101 DeepLab 43.2 58.5 93.5 84.0 15.1 54.6
Ensemble C 73.9 47.4 99.1 92.5 52.7 34.7
Ensemble A 73.8 47.0 99.1 92.6 38.6 36.9
PixOOD w/ DeepLab Decoder 69.3 23.5 97.4 76.0 13.8 69.9
DeiT III (IN1K), ViT-S, 16x16 patch size, linear decoder 44.9 54.7 97.9 89.2 1.7 54.2
PixOOD 64.8 22.4 96.3 73.1 17.6 76.6
SegFormer-B5 67.4 24.1 97.2 77.2 30.7 71.9
ObsNet-R101-DLv3plus 65.3 32.1 98.5 87.8 45.6 63.4
Mask2Former-SwinB 67.2 13.2 90.4 47.0 55.1 82.8
Physically Feasible Semantic Segmentation 67.4 42.2 98.7 89.4 4.8 48.3
@@
Another noteworthy observation is that all our models have relatively low ECE values, indicating that they are well-calibrated, even though no explicit calibration techniques were applied. Even the DeiT-III-based model, which scores low on the overall BRAVO score, achieves a low ECE value of 1.7. Therefore, further investigation is needed to understand why the ECE values are so low.
@@
Finally, while scaling from ViT-L to ViT-g improves mIoU for valid pixels (see Table 3), OOD detection performance shows a noticeable degradation.
@@ 5 Conclusion
This report presents the winning solution for Track 1 of the 2024 BRAVO Challenge. Our approach is based on fine-tuning Vision Foundation Models (VFMs) for semantic segmentation, leveraging the strong representations learned during pre-training. The results demonstrate that, just through better pre-training, a fine-tuned VFM is more robust under distribution shifts than complex specialist models, and can both accurately classify pixels and reliably estimate its confidence in these predictions. We evaluated our approach in several configurations, which has lead to several new observations. While we identify potential causes of these observations, future work is necessary to explore this further. Particularly, we believe it would be interesting to investigate the reason for the difference in semantic and OOD metrics with the Mask2Former decoder compared to the linear decoder. Additionally, future works could explore whether some of the more specialized out-of-distribution detection or calibration methods are still effective when used in combination with VFMs.